# Assessing the accuracy of the star formation rate measurements by direct star count in molecular clouds

### v2

@@
Assessing the accuracy of the star formation rate measurements by direct star count in molecular clouds
Star formation estimates based on the counting of Young Stellar Objects (YSOs) is commonly applied to nearby star-forming regions in the Galaxy and, in principle, could be extended to any star-forming region where direct star counts are possible. With this method, the SFRs are measured using the counts of YSOs in a particular Class, a typical mass, and the lifetime associated with this Class. Another variant of this method is to use the total number of YSOs found in a star-forming region along with a characteristic YSO timescale. However, the assumptions underlying the validity of the method such as that of a constant star formation history (SFH) and whether the method is valid for all protostellar Classes has never been fully tested. In this work, we use Monte Carlo models to test the validity and robustness of the method. We build synthetic clusters in which stars form at times that are randomly drawn from a specified SFH distribution function. The latter is either constant or time-dependent with a burst like behavior. The masses of the YSOs are randomly drawn from a stellar initial mass function (IMF) which can be either similar to that of the Milky Way field or be variable within the limits of variations observed among young stellar clusters in the Galaxy. For each star in every cluster, the lifetimes associated with the different protostellar classes are also randomly drawn from Gaussian distribution functions centered around their most likely value as suggested by the observations. We find that only the SFR derived using the Class 0 population can reproduce the true SFR at all epochs, and this is true irrespective of the shape of the SFH. For a constant SFH, the SFR derived using the more evolved populations of YSOs (Class I, Class F, Class II, and Class III) reproduce the real SFR only at later epochs which correspond to epochs at which their numbers have reached a steady state. For a time-dependent burst-like SFH, all SFR estimates based on the number counts of the evolved populations fail to reproduce the true SFR. We show that these conclusions are independent of the IMF. We argue that the SFR based on the Class 0 alone can yield reliable estimates of the SFR. We also show how the offsets between Class I and Class II based SFRs and the true SFR plotted as a function of the number ratios of Class I and Class II versus Class III YSOs can be used in order to provide information on the SFH of observed molecular clouds.
@@
Understanding the rate at which molecular clouds convert their gas reservoirs into stars is of particular importance for many branches of Astrophysics. The star formation rate (SFR) determines a wide range of the observed properties of galaxies and affects their chemical and dynamical evolution (Helou et al. 2000; Dib et al. 2009; Maraston et al. 2010; Dib 2011; Dariush et al. 2016; Pacifici et al. 2016; Barrera-Ballesteros et al. 2021; Grisdale et al. 2022; Garduño et al. 2023; Egorov et al. 2023). Within individual molecular clouds, the SFR is regulated by a number of physical processes that include the clouds’ self-gravity, supersonic turbulence and the way it is driven, chemical composition, the incident radiation field, and tidal fields (Krumholz & McKee 2005; Dib et al. 2007,2008,2010; Padoan & Nordlund 2011; Glover & Clark 2012; Federrath & Klessen 2012; Dib et al. 2020; Schneider et al. 2022; Rani et al. 2022; Dib 2023a, Li 2024; Zhou et al. 2024a,b). On large scales, the SFR can be regulated by other processes such as the gravity due to the existing stellar populations (Dib et al. 2017a; Shi et al. 2018; Marchuk 2018; Pessa et al. 2021), galactic shear (Seigar et al. 2005; Dib et al. 2012; Aouad et al. 2020), and the overall galactic potential (Jog 2014; Meidt et al. 2018).
@@
The SFR can be measured using a variety of star formation indicators. Some of these methods rely on the stellar light that is directly emitted by young stars such as in the ultraviolet, or by dust processed stellar light in the infrared. For a detail review on these methods, we refer the reader to Calzetti (2013). While it should be noted that these methods are mostly employed for external galaxies where stellar populations are not well resolved, they make assumptions about the shape of the stellar initial mass function, and commonly assume that it universal within and across galaxies. Recent work has shown that this is not the case, neither on individual molecular clouds scales (Dib 2014; Dib et al. 2017b; Dib 2023b) nor for entire galaxies (Dib & Basu 2018; Dib 2022). In nearby star forming regions in the Milky Way, a different method is usually used to infer the SFR. This method is based on the direct counting of young stellar objects (YSOs)111We use the acronym YSO to designate the young stellar population in its ensemble. When referring specifically to young YSOs (Class 0, Class I, and the F Class), we employ the terms protostars. The older YSOs (Class II and Class III), are referred to as pre-main sequence (PMS) stars. which are optimally identified in the infrared (Forbrich 2009; Enoch et al. 2009; Evans et al. 2009; Heiderman et al. 2010; Guthermuth et al. 2011; Ybarra et al. 2013; Jose et al. 2013; Strafella et al. 2015; Stutz & Kainulainen 2015; Jose et al. 2016; Fischer et al. 2016; Hasan et al. 2023; Tobin & Sheehan 2024). The number of YSOs in a given evolutionary stage in a star-forming region, is converted into a using
@@
where the index runs over the different protostellar classes (=0, I, F, II, and III for objects in Class 0, Class I, Class F, Class II, and Class III, respectively), is the mean protostellar mass present in the region, and is the YSOs typical lifetime in each class. The accuracy of the SFR measurements using Eq. 1 depends on the validity of some of the assumptions that are adopted in calculating the average mass and protostellar lifetimes. For example, for Class II YSOs, is often taken to be Myr. The mean mass is usually assumed to be M⊙ which corresponds to the mean mass in a Milky-Way like IMF. These types of simplifications could potentially generate a bias in the estimates of the SFR. Another source of uncertainty that can affect the SFR measurements using YSO counts is the incompleteness of the YSO censuses and the level of incompleteness can vary with the evolutionary stage of the YSOs as well as with the specifics of how the observational data was obtained and reduced (Gutermuth et al. 2009; Winston et al. 2010; Broekhoven-Fiene et al. 2014; Dunham et al. 2015; Gutermuth & Heyer 2015). In this work, we attempt to estimate if such biases exist and what are their amplitudes. We approach this problem by generating synthetic populations of YSOs that form in a molecular cloud with a prescribed star formation history. The YSOs have masses that are drawn from an IMF and are assigned protostallar lifetimes that span the realistic range of each class. We then measure the real star formation rate and compare it to the one that is derived using Eq. 1 for YSOs in different classes. In §. 2, we describe the elements of the model, namely the protostellar mass function from which stellar masses are drawn, the function that describes the star formation history (SFH) of the YSOs, and their lifetimes associated with in each Class. In §. 3, we present the results for the time evolution of the SFRs derived from the various protostellar Classes and compare them to the real SFR, exploring both the effects of the SFH and of the IMF. Finally in §. 4 we measure the offsets between the real SFRs and those derived using the Class I and Class II population. We explore the correlation between these offsets and the ratios between the numbers of Class I and II YSOs and those of the Class 0. In §. 5, we conclude.
@@ 2.1 The protostellar mass function
Since we do not account for the mass evolution of YSOs, which is minimal for low mass stars, we also refer to the protostellar mass function as the stellar initial mass function (IMF) and assume that it can be described with a tapered power-law function (TPL; Dib 2014; Dib et al. 2017b; Dib 2023b). The TPL is characterized by two power-laws in the low and high-mass regimes, and a characteristic mass (i.e., peak mass) and is given by
where is a normalization term, is the exponent of the power-law in the high mass end, is the exponent in the low-mass regime, and is the characteristic mass. The Salpeter (1955) value for the slope in the high-mass regime corresponds to . The Milky Way values for the two other variables for a single-star IMF (i.e., corrected for binarity) are and M⊙ (Parravano et al. 2011). In this work, we consider cases where the parameters of the clusters’ IMF are assigned Milky Way-like values, and others where we vary the three parameters within the ranges that are permitted by the observations of young Galactic stellar clusters. The minimum and maximum stellar masses we consider are 0.02 M⊙ and 50 M⊙, respectively.
@@ 2.2 Protostellar classes and associated lifetimes
The classification into Classes I, II, and III was initially proposed by Lada & Wilking (1984) and is based on the value of the slope of the spectral energy distribution (SED) in the wavelength range of 2 to 20 m. It was subsequently used in numerous studies (André & Montmerle 1994; Evans et al. 2009; Rebull et al. 2014; Furlan et al. 2016; Sharma et al. 2017; Pokhrel et al. 2020; Kuhn et al. 2021; Sun et al. 2022). The Class 0 classification was later introduced by André et al. (2000) and is based on the indirect inference of the presence of a YSO based on the detection of a compact submillimeter source and a collimated CO outflow. These protostellar classes are associated with different phases of the formation of stars and each phase is characterized by a specific shape of the spectral energy distribution (SED). These phases correspond to: in Class 0, the formation of a YSO in the central region of a protostellar core with an envelope mass that is much in excess of the YSO mass, for Class I: the collapse of the envelope onto the central object with the transition between Class 0 and Class I being the point in time at which the envelope mass and the mass of the protostar are nearly equal, for Class II: the emergence of a disk around the central star, and for Class III: the dissipation of the disk by various processes such as the formation of planets, photo-evaporation, and tidal stripping. An intermediate class between Class 0 and Class I has been proposed by Greene et al. (1994) and was labelled the Flat class (i.e., a relatively flat SED in the wavelength range of 2 to 20 m) but it was shown that this class is close to Class I in terms of the evolutionary status of the YSOs as it is associated with signs of a collapsing envelope (Calvet et al. 1994). Recent work shows that genuine flat-spectrum YSOs are rare, once contamination by reddened Class II YSOs and dusty star-forming galaxies is corrected for (Pokhrel et al. 2023)222The exact nature of the flat-spectrum YSOs is still under investigations. A recent analysis using observations from the ATACAMA Compact Array (ACA) suggests that these objects are more likely disk-dominated than envelope dominated and, as such, are more likely to be late-stage protostars or edge-on Class II YSOs (Federman et al. 2023).
@@
The exact duration of each evolutionary stage for YSOs classes is relatively uncertain. Estimates of the mean duration of each of these stages depend on the number of objects found in each class in star-forming regions and which may be affected by misclassifications due for example to YSOs being seen edge-on (Gutermuth et al. 2009). The usual approach is to use Eq. 1 and the numbers of YSOs found in different classes in order to derive relative ages for each class. The assumption is made that YSOs are formed at a constant rate over time and that their formation occurs over timescales that are typically longer than the longest duration of any class. By constraining the age of one of the classes such as that of the Class II T Tauri stars, which is assumed to be of the order of their disk lifetime 333The disk lifetime of a class II YSO is estimated as being the timescale over which the fraction of YSOs with a disk decays as a function of the mean isochronal ages of star forming regions of a certain range of ages (Haish et al. 2001; Hernández et al. 2007,2008; Mamajek 2009)., one can then derive the corresponding duration of other classes for each YSO. A number of studies measured these timescales in nearby star-forming clouds and found them to fall in the range to Myr for Class 0 YSOs (André et al. 1993; Froebrich et al. 2006; Enoch et al. 2009; Fischer et al. 2017; Kristensen & Dunham 2018), between and for Class I YSOs (Wilking et al. 1989; Greene et al. 1994; Kenyon & Hartmann 1995; Hatchell et al. 2007), and between to Myr for Class II YSOs (Wilking et al. 1989; Kenyon & Hartmann 1995; Hernández et al. 2008) and ages in excess of 2 Myr for Class III YSOs (Wilking et al. 2005).
@@
In this work, we adopt the values of the lifetimes measured by Evans et al. (2009). These were derived from observations of several nearby molecular clouds in the the Spitzer legacy project, ’From Molecular Cores to Planet-forming Disks’ (c2d). Evans et al. (2009) found that the mean lifetimes of Class 0, Class I and the Flat Class are Myr, Myr and Myr. These were derived assuming a constant SFR and a lifetime for Class II of Myr and include a correction of the photometry for extinction effects. The uncertainties on these lifetimes are quite large. For the Class II, the uncertainty is about Myr. Given all the assumptions such as that of a constant SFR, the potential confusion of a fraction of the YSOs with extragalactic sources, and differences in the relative numbers of YSOs found in each region of the c2d cloud sample, we adopt uncertainties on the lifetimes that are of their adopted values, such that Myr, Myr, and Myr. These results were corroborated by the study of Hsieh & Lai (2013) who re-analyzed the c2d data and found similar results for the lifetimes than those obtained by Evans et al. (2009). For the Class III YSOs, their census in the c2d is incomplete and therefore, their associated lifetimes were not estimated. In general, the lifetime associated with Class III YSOs is uncertain and a ”reasonable” estimate for it would the duration it takes a young star to descend the Hayashi tracks and reach the main sequence after the disk has dispersed. An approximation of this duration is the Kelvin-Helmotz times scale which is the typical time over which a young star contracts and radiates its kinetic energy and descends the Hayashi track to reach the main sequence (Hayashi 1961). The Kelvin-Helmotz timescale is given by , where is the gravitational constant and , , and are the star’s mass, radius, and stellar luminosity, respectively. For a solar mass stars, Myrs and it is lower for higher mass stars and substantially lower for low mass stars. However, as we will be following the evolution of the young stellar populations for the first Myrs, the choice of any value of that is larger than 30 Myrs will make little difference to our calculations. Here, we consider a single value across the entire stellar mass range of Myr with Myr.
@@
In this work, we draw the lifetimes of the population of YSOs from Gaussian distribution functions given by:

### v4

# Assessing the accuracy of star formation rate measurements by direct star count in molecular clouds

@@
Assessing the accuracy of star formation rate measurements by direct star count in molecular clouds
Star formation estimates based on the counting of young stellar objects (YSOs) are commonly carried out for nearby star-forming regions in the Galaxy, and in principle could be extended to any star-forming region where direct star counts are possible. With this method, the SFRs are measured using the counts of YSOs in a particular class, a typical mass, and the lifetime associated with this class. Another variant of this method is to use the total number of YSOs found in a star-forming region along with a characteristic YSO timescale. However, the assumptions underlying the validity of this method, such as that of a constant star formation history (SFH), have never been fully tested, and it remains unclear as to whether or not the method is valid for all protostellar classes. In this work, we use Monte Carlo models to test the validity and robustness of the method. We build synthetic clusters in which stars form at times that are randomly drawn from a specified SFH distribution function. The latter is either constant or time dependent, with a burst like behavior. The masses of the YSOs are randomly drawn from a stellar initial mass function (IMF), which can be either similar to that of the Milky Way field or be variable within the limits of the variations observed among young stellar clusters in the Galaxy. For each star in every cluster, the lifetimes associated with the different protostellar classes are also randomly drawn from Gaussian distribution functions centered around their most likely value as suggested by the observations. We find that only the SFR derived using the Class 0 population can reproduce the true SFR at all epochs, and this is true irrespective of the shape of the SFH. For a constant SFH, the SFR derived using the more evolved populations of YSOs (Class I, Class F, Class II, and Class III) reproduce the real SFR only at later epochs, which correspond to epochs at which their numbers have reached a steady state. For a time-dependent burst-like SFH, all SFR estimates based on the number counts of the evolved populations fail to reproduce the true SFR. We show that these conclusions are independent of the IMF. We argue that the SFR based on the Class 0 alone can yield reliable estimates of the SFR. We also show how the offsets between Class I- and Class II-based SFRs and the true SFR plotted as a function of the number ratios of Class I and Class II versus Class III YSOs can be used in order to provide information on the SFH of observed molecular clouds.
@@
Understanding the rate at which molecular clouds convert their gas reservoirs into stars is of particular importance for many branches of astrophysics. The star formation rate (SFR) determines a wide range of the observed properties of galaxies and affects their chemical and dynamical evolution (Helou et al. 2000; Dib et al. 2009; Maraston et al. 2010; Dib 2011; Dariush et al. 2016; Pacifici et al. 2016; Barrera-Ballesteros et al. 2021; Grisdale et al. 2022; Garduño et al. 2023; Egorov et al. 2023). Within individual molecular clouds, the SFR is regulated by a number of physical processes that include the clouds’ self-gravity, supersonic turbulence and the way it is driven, their chemical composition, the incident radiation field, and tidal fields (Krumholz & McKee 2005; Dib et al. 2007,2008,2010; Padoan & Nordlund 2011; Glover & Clark 2012; Federrath & Klessen 2012; Dib et al. 2020; Schneider et al. 2022; Rani et al. 2022; Dib 2023a, Li 2024; Zhou et al. 2024a,b). On large scales, the SFR can be regulated by other processes, such as the gravity due to the existing stellar populations (Dib et al. 2017a; Shi et al. 2018; Marchuk 2018; Pessa et al. 2022), galactic shear (Seigar et al. 2005; Dib et al. 2012; Aouad et al. 2020), and the overall galactic potential (Jog 2014; Meidt et al. 2018).
@@
The SFR can be measured using a variety of star-formation indicators. Some of these methods rely on the stellar light that is directly emitted by young stars, such as in the ultraviolet, or by dust-processed stellar light in the infrared. For a detail review of these methods, we refer the reader to Calzetti (2013). While it should be noted that these methods are mostly employed for external galaxies where stellar populations are not well resolved, they make assumptions about the shape of the stellar IMF, and commonly assume that it is universal within and across galaxies. Recent work has shown that this is not the case, either on individual molecular clouds scales (Dib 2014; Dib et al. 2017b; Dib 2023b) or for entire galaxies (Dib & Basu 2018; Dib 2022). In nearby star forming regions in the Milky Way, a different method is usually used to infer the SFR. This method is based on the direct counting of young stellar objects (YSOs)111We use the acronym YSO to designate the young stellar population in its ensemble. When referring specifically to young YSOs (Class 0, Class I, and the F Class), we employ the terms protostars. The older YSOs (Class II and Class III) are referred to as pre-main sequence (PMS) stars., which are optimally identified in the infrared (Forbrich 2009; Enoch et al. 2009; Evans et al. 2009; Heiderman et al. 2010; Guthermuth et al. 2011; Ybarra et al. 2013; Jose et al. 2013; Strafella et al. 2015; Stutz & Kainulainen 2015; Jose et al. 2016; Fischer et al. 2016; Hasan et al. 2023; Tobin & Sheehan 2024). The number of YSOs in a given evolutionary stage in a star-forming region, , is converted into a using
@@
where the index runs over the different protostellar classes (=0, I, F, II, and III for objects in Class 0, Class I, Class F, Class II, and Class III, respectively), is the mean protostellar mass present in the region, and is the typical lifetime of the YSOs spent in each class. The accuracy of the SFR measurements using Eq. 1 depends on the validity of some of the assumptions that are adopted when calculating the average mass and protostellar lifetimes. For example, for Class II YSOs, is often taken to be Myr. The mean mass is usually assumed to be M⊙, which corresponds to the mean mass in a Milky Way-like IMF. These types of simplifications could potentially generate a bias in the estimates of the SFR. Another source of uncertainty that can affect the SFR measurements using YSO counts is the incompleteness of the YSO censuses and the level of incompleteness can vary with the evolutionary stage of the YSOs as well as with the specific details of how the observational data were obtained and reduced (Gutermuth et al. 2009; Winston et al. 2010; Broekhoven-Fiene et al. 2014; Dunham et al. 2015; Gutermuth & Heyer 2015). In this work, we attempt to estimate whether or not such biases exist and to determine their amplitudes. We approach this problem by generating synthetic populations of YSOs that form in a molecular cloud with a prescribed star formation history (SFH). The YSOs have masses that are drawn from an IMF and are assigned protostellar lifetimes that span the realistic range of each class. We then measure the real SFR and compare it to the SFR derived using Eq. 1 for YSOs in different classes. In §. 2, we describe the elements of the model, namely the protostellar mass function from which stellar masses are drawn, the function that describes the SFH of the YSOs, and their lifetimes spent within each class. In §. 3, we present the results of the time evolution of the SFRs derived from the various protostellar classes and compare them to the real SFR, exploring both the effects of the SFH and of the IMF. In §. 4 we measure the offsets between the real SFRs and those derived using the Class I and Class II populations. We explore the correlation between these offsets and the ratios between the numbers of Class I and II YSOs and those of Class III. Finally, in §. 5, we present our conclusions.
@@ 2.1 The protostellar mass function
As we do not account for the mass evolution of YSOs, which is minimal for low-mass stars, we also refer to the protostellar mass function as the stellar IMF and assume that it can be described with a tapered power-law function (TPL; Dib 2014; Dib et al. 2017b; Dib 2023b). The TPL is characterized by two power laws in the low- and high-mass regimes, and a characteristic mass (i.e., peak mass) and is given by
where is a normalization term, is the exponent of the power law in the high-mass end, is the exponent in the low-mass regime, and is the characteristic mass. The Salpeter (1955) value for the slope in the high-mass regime corresponds to . The Milky Way values for the two other variables for a single-star IMF (i.e., corrected for binarity) are and M⊙ (Parravano et al. 2011). In this work, we consider cases where the parameters of the clusters’ IMF are assigned Milky Way-like values, and others where we vary the three parameters within the ranges that are permitted by the observations of young Galactic stellar clusters. The minimum and maximum stellar masses we consider are 0.02 M⊙ and 50 M⊙, respectively.
@@ 2.2 Protostellar classes and associated lifetimes
The classification into Classes I, II, and III was initially proposed by Lada & Wilking (1984) and is based on the value of the slope of the spectral energy distribution (SED) in the wavelength range of 2 to 20 m. This classification system was subsequently used in numerous studies (André & Montmerle 1994; Evans et al. 2009; Rebull et al. 2014; Furlan et al. 2016; Sharma et al. 2017; Pokhrel et al. 2020; Kuhn et al. 2021; Sun et al. 2022). The Class 0 classification was later introduced by André et al. (2000) and is based on the indirect inference of the presence of a YSO based on the detection of a compact submillimeter source and a collimated CO outflow. These protostellar classes are associated with different phases of the formation of stars and each phase is characterized by a specific shape of the SED. The star-formation processes taking place in these phases are as follows: In Class 0, a YSO is formed in the central region of a protostellar core with an envelope mass that is far in excess of the YSO mass; in Class I, there is collapse of the envelope onto the central object with the transition between Class 0 and Class I being the point in time at which the envelope mass and the mass of the protostar are almost equal; in Class II, a disk emerges around the central star; and in Class III, the disk dissipates by various processes, such as the formation of planets, photo-evaporation, and tidal stripping. An intermediate class between Class 0 and Class I has been proposed by Greene et al. (1994) and was labelled the Flat class (i.e., a relatively flat SED in the wavelength range of 2 to 20 m) but it was shown that this class is close to Class I in terms of the evolutionary status of the YSOs as it is associated with signs of a collapsing envelope (Calvet et al. 1994). Recent work shows that genuine flat-spectrum YSOs are rare, once contamination by reddened Class II YSOs and dusty star-forming galaxies is corrected for (Pokhrel et al. 2023)222The exact nature of the flat-spectrum YSOs is still under investigation. A recent analysis using observations from the ATACAMA Compact Array (ACA) suggests that these objects are more likely disk dominated than envelope dominated and, as such, are more likely to be late-stage protostars or edge-on Class II YSOs (Federman et al. 2023).
@@
The exact duration of each evolutionary stage for YSO classes is relatively uncertain. Estimates of the mean duration of each of these stages depend on the number of objects found in each class in star-forming regions and this may be affected by misclassifications due for example to YSOs being seen edge-on (Gutermuth et al. 2009). The usual approach is to use Eq. 1 and the numbers of YSOs found in different classes in order to derive relative ages for each class. The assumption is made that YSOs are formed at a constant rate over time and that their formation occurs over timescales that are typically longer than the longest duration of any class. By constraining the age of one of the classes, such as that of the Class II T Tauri stars, which is assumed to be of the order of their disk lifetime 333The disk lifetime of a class II YSO is estimated as being the timescale over which the fraction of YSOs with a disk decays as a function of the mean isochronal ages of star forming regions of a certain range of ages (Haish et al. 2001; Hernández et al. 2007,2008; Mamajek 2009)., one can then derive the corresponding duration of the other classes for each YSO. A number of studies measured these timescales in nearby star-forming clouds and found them to fall in the range of to Myr for Class 0 YSOs (André et al. 1993; Froebrich et al. 2006; Enoch et al. 2009; Fischer et al. 2017; Kristensen & Dunham 2018), between and for Class I YSOs (Wilking et al. 1989; Greene et al. 1994; Kenyon & Hartmann 1995; Hatchell et al. 2007), and between and Myr for Class II YSOs (Wilking et al. 1989; Kenyon & Hartmann 1995; Hernández et al. 2008), and found ages in excess of 2 Myr for Class III YSOs (Wilking et al. 2005).
@@
In the present work, we adopt the values of the lifetimes measured by Evans et al. (2009). These were derived from observations of several nearby molecular clouds in the the Spitzer legacy project, “From Molecular Cores to Planet-forming Disks” (c2d). Evans et al. (2009) found that the mean lifetimes of Class 0, Class I, and the Flat Class are Myr, Myr, and Myr. These were derived assuming a constant SFR and a lifetime for Class II of Myr and include a correction of the photometry for extinction effects. The uncertainties on these lifetimes are quite large. For Class II, the uncertainty is about Myr. Given all the assumptions, such as that of a constant SFR, the potential confusion of a fraction of the YSOs with extragalactic sources, and differences in the relative numbers of YSOs found in each region of the c2d cloud sample, we adopt uncertainties on the lifetimes that are of their adopted values, such that Myr, Myr, and Myr. These results were corroborated by the study of Hsieh & Lai (2013), who reanalyzed the c2d data and found similar results for the lifetimes to those obtained by Evans et al. (2009). For the Class III YSOs, their census in the c2d is incomplete, and therefore their associated lifetimes were not estimated. In general, the lifetime associated with Class III YSOs is uncertain. An approximation of this duration is the Kelvin-Helmotz timescale, which is the typical time during which a young star contracts and radiates its kinetic energy and descends the Hayashi track to reach the main sequence (Hayashi 1961). The Kelvin-Helmotz timescale is given by , where is the gravitational constant and , , and are the star’s mass, radius, and stellar luminosity, respectively. For solar-mass stars, Myr, and this value is lower for higher-mass stars and is substantially higher for low-mass stars. However, as we follow the evolution of the young stellar populations for the first Myr, the choice of any value of that is larger than 30 Myr will make little difference to our calculations. Here, we consider a single value across the entire stellar mass range of Myr with Myr.
@@
In the present work, we draw the lifetimes of the population of YSOs from Gaussian distribution functions given by